OpenStudy (anonymous):

use factoring to solve this equation x^2-x=6

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

Before you can factor, you need to subtract 6 from both sides. \[x^{2}-x-6=0\]

OpenStudy (anonymous):

Now you want two numbers that multiply to 6 but are different by 1. Can you think of two numbers that work for this?

OpenStudy (anonymous):

2 and 3

OpenStudy (anonymous):

x=3 is a root. (x^2-x-6)/(x-3)=x+2 Thus: x^2-x-6 = (x-3)(x+2) = 0 x1 = 3 and x2 = -2

OpenStudy (anonymous):

Very good. So let's rewrite the middle term\[x^{2}-3x+2x-6=0\]and then regroup\[(x^{2}-3x) +(2x-6)\]

OpenStudy (anonymous):

We can factor an x out of the first group and a 2 out of the second and then we'll see something nice

OpenStudy (anonymous):

\[x(x-3)+2(x-3)\]Now, factor out the x-3 from both sets\[(x-3)(x+2)\]and we're done!